Form a polynomial f(x) with real coefficients having the given degree and zeros; degree: 4; zeros: 3-5i, 5; multiplicity: 2
Answer by ewatrrr(24785)

f(x) = (3-5i)(3+5i)(x-5)^2 = (9 + 25)(x-5)^2 = 34(x-5)^2 Expand If You wish